### v1

# LLM-Based Multi-Hop Question Answering with Knowledge Graph Integration in Evolving Environments

Abstract
The rapid obsolescence of information in Large Language Models (LLMs) has driven the development of various techniques to incorporate new facts. However, existing methods for knowledge editing still face difficulties with multi-hop questions that require accurate fact identification and sequential logical reasoning, particularly among numerous fact updates. To tackle these challenges, this paper introduces Graph Memory-based Editing for Large Language Models (GMeLLo), a straitforward and effective method that merges the explicit knowledge representation of Knowledge Graphs (KGs) with the linguistic flexibility of LLMs. Beyond merely leveraging LLMs for question answering, GMeLLo employs these models to convert free-form language into structured queries and fact triples, facilitating seamless interaction with KGs for rapid updates and precise multi-hop reasoning. Our results show that GMeLLo significantly surpasses current state-of-the-art knowledge editing methods in the multi-hop question answering benchmark, MQuAKE, especially in scenarios with extensive knowledge edits.
[…]
[ linecolor=black, leftmargin=0, rightmargin=0, backgroundcolor=gray!10, innertopmargin=5pt, innerbottommargin=5pt, skipabove=skipbelow=]examplebox
LLM-Based Multi-Hop Question Answering with Knowledge Graph Integration in Evolving Environments
1 Introduction
Figure 1: Multi-hop question answering in dynamic domains Zhong et al. (2023). Dynamic nature of information: Changes over time may trigger subsequent modifications. For instance, a transition in the British Prime Minister, such as from Boris Johnson to Rishi Sunak, necessitates corresponding adjustments, like the change in the British Prime Minister’s spouse.
As the widespread deployment of LLMs continues, the imperative to keep their knowledge accurate and up-to-date, without incurring extensive retraining costs, becomes increasingly evident Sinitsin et al. (2020). Several approaches have been proposed in prior works to address this challenge, with some focusing on the incremental injection of new facts into language models Rawat et al. (2020); De Cao et al. (2021); Meng et al. (2022); Mitchell et al. (2022a). Interestingly, certain methodologies in the literature diverge from the conventional path of updating model weights, opting instead for an innovative strategy involving the use of external memory to store the edits Mitchell et al. (2022b); Zhong et al. (2023).
As LLMs operate as black boxes, modifying one fact might inadvertently alter another, making it challenging to guarantee accurate revisions. In this paper, we introduce GMeLLo, an effective approach designed to synergize the strengths of LLMs and KGs in addressing the multi-hop question answering task after knowledge editing Zhong et al. (2023). An illustrative example of our focus is presented in Figure  1. Following an update regarding the information of the British Prime Minister, it becomes evident that the corresponding spouse information should also be modified.
As depicted in Figure  2, our GMeLLo comprises the following key steps:
  • We utilize LLMs to translate edited fact sentences into triples, employing these triples to update the KG and ensure its information remains up to date.
  • Given a question, we utilize LLMs to extract its relation chain, encompassing the primary entity and its connections with other unknown entities. After populating a template, we convert the relation chain into a formal query and use it to search the updated KG.
  • In addition, we retrieve the most pertinent edited facts based on the question and prompt LLMs to generate an answer in accordance with these facts.
  • In instances where the answer provided by the LLM conflicts with that from the KG, we prioritize the answer from the KG as the final response.
LLMs, trained on extensive sentence corpora Brown et al. (2020); Rae et al. (2022); Chowdhery et al. (2023), are expected to encapsulate a wide range of commonly used sentence structures. As a result, they are invaluable tools for analyzing sentences and extracting entities and relations. Once the correct relation chain and edited triples are obtained, using a formal query to interrogate the KG in a Knowledge-based Question Answering (KBQA) Cui et al. (2017) manner ensures precision in the searching process. In cases where KBQA fails, we still have LLMs for question answering (QA) to ensure comprehensive coverage. GMeLLo outperforms current state-of-the-art (SOTA) methods on two datasets from the MQuAKE benchmark, affirming its effectiveness in multi-hop question answering within an evolving environment.
2 Related Work
This work utilizes both KGs and LLMs to address the challenge of multi-hop question answering, with a particular focus on scenarios involving evolving factual knowledge. Therefore, we review existing literature on multi-hop question answering, knowledge editing, and the augmentation of LLMs with knowledge graphs [^1] Due to space constraints, some of the literature is located in Appendix B..
2.1 Multi-Hop Question Answering
Multi-hop question answering is more challenging because it requires not only recalling facts but also appropriately aggregating and chaining them. Facts can be sourced from a knowledge graph (Lin et al., 2018; Cheng et al., 2023; Zhong et al., 2023), tables (Yin et al., 2016), free-form text (Yang et al., 2018; Welbl et al., 2018), or a heterogeneous combination of these sources (Chen et al., 2020; Mavi et al., 2022; Lei et al., 2023). With the development of LLMs, prompt-based methods combined with an optional retrieval module have become a popular approach for handling multi-hop question answering Khattab et al. (2022); Press et al. (2023); Zhong et al. (2023). While most previous works focus on a static information base, our approach targets a dynamic domain, accommodating changes in facts.
2.2 Knowledge Editing
As highlighted in Yao et al. (2023), two paradigms exist for editing knowledge: modifying model parameters and preserving model parameters.
2.2.1 Modifying Model Parameters
In the case of modifying model parameters, this can be further categorized into meta-learning or locate-and-edit approaches. Meta-learning methods, as discussed in De Cao et al. (2021); Mitchell et al. (2022a), utilize a hyper network to learn the necessary adjustments for editing LLMs. The locate-then-edit paradigm, as demonstrated in Dai et al. (2022); Meng et al. (2022, 2023); Li et al. (2023a); Gupta et al. (2023); Zhang et al. (2024), involves initially identifying parameters corresponding to specific knowledge and subsequently modifying them through direct updates to the target parameters.
2.2.2 Preserving Model Parameters
In the case of preserving model parameters, the introduction of additional parameters or external memory becomes necessary. The paradigm of additional parameters, as presented in Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2022), incorporates extra trainable parameters into the language model. These parameters are trained on a modified knowledge dataset, while the original model parameters remain static. On the other hand, memory-based models Mitchell et al. (2022b); Zhong et al. (2023) explicitly store all edited examples in memory and employ a retriever to extract the relevant edit facts for each new input, guiding the model in generating the edited output.
While previous evaluation paradigms have primarily focused on validating the recall of edited facts, Zhong et al. (2023) introduced MQuAKE, a benchmark that includes multi-hop questions involving counterfactual or temporal edits. The two datasets within MQuAKE assess whether methods can accurately answer questions where the response should change due to edited facts.
Figure 2: The illustration depicts our proposed method, GMeLLo. We begin by utilizing LLMs to extract entities and relations from edited facts, resulting in a list of edited fact triples. These triples are then used to update a KG. Similarly, we employ LLMs to extract relation chains from a given question. By populating this information into a template, we generate a formal query suitable for use in KBQA Lan et al. (2022). Simultaneously, we utilize LLMs for question answering, providing an answer based on the relevant edited facts retrieved. In cases where the LLM’s answer contradicts that of the KG, we defer to the KG’s answer as the final response.
3 GMeLLo: Graph Memory-based Editing for Large Language Models
In this section, we introduce our method GMeLLo for multi-hop question answering with knowledge editing (Figure 2).
3.1 Extracting Fact Triples from Edited Information Using LLMs
KGs play a pivotal role in enhancing the capabilities of LLMs by offering external knowledge for improved inference and interpretability, as demonstrated by recent studies Pan et al. (2023); Rawte et al. (2023). Apart from merely storing updated information in an external memory, such as a list of separate sentence statements as seen in conventional approaches Zhong et al. (2023), we utilize the KG to maintain inherent connections and ensure the integration of the latest information.
In our approach, we leverage Wikidata Vrandečić and Krötzsch (2014), a widely recognized KG, as the foundational knowledge base. When updated facts are received, we utilize LLMs to extract entities from the sentences and determine their relationships (selecting a relation from the predefined list). This process generates edited fact triples, which are then used to update the KG (see Figure  2). Updating the KG with an edited fact triple involves identifying the connections in the KG based on the subject entity and relation, breaking these connections, and establishing a new connection based on the triple.
We incorporate in-context learning Dong et al. (2023) to ensure the LLMs have thorough understanding of the task. Furthermore, given the possibility that LLMs may generate relations not present in the predefined relation list Chen et al. (2024), we use a retrieval model to identify the most similar relation (i.e., the closest relation in the embedding space) from the predefined relation list. The integration of retrieval model makes the triple extraction process more robust.
3.2 Extracting Relation Chain from Questions Using LLMs
As the world evolves rapidly, the training data for LLMs can quickly become outdated. However, since the evolution of linguistic patterns typically progresses at a slower pace, the extensive training data of LLMs should enable them to effectively comprehend most sentence patterns. In this paper, we employ LLMs to extract the relation chain from a sentence, encompassing the mentioned entity in the question and its relations with other unidentified entities. Similar to the fact triple exaction 3.1, we task LLMs with selecting a relation from a predefined list to mitigate varied representations of the same relation. Take a question sentence from the MQuAKE-CF Zhong et al. (2023) dataset as an example,
{examplebox}
Question
What is the capital of the country of citizenship of the child of the creator of Eeyore?
Relation Chain
Eeyore->creator->?x->child->?y
->country of citizenship
->?z->capital->?m
The presented question necessitates a 4-hop reasoning process. With "Eeyore" as the known entity in focus, the journey to the final answer involves identifying its creator ’?x’, moving on to the creator’s child ’?y’, obtaining the child’s country of citizenship ’?z’, and culminating with the retrieval of the country’s capital ’?m’. All the relations, such as ’creator,’ ’child,’ ’country of citizenship,’ and ’capital,’ are chosen from a predefined list of relations. The relation chain encapsulates all essential information for deriving the answer.
To enable LLMs to extract relation chains and generate outputs in a structured template, we provide several examples of relation chain extraction in the prompt and utilize in-context learningDong et al. (2023), as detailed in Appendix A.4.
3.3 Converting Relation Chain into a Formal Query
Once the relation chain is obtained, the next step involves integrating the known entity and the relations into a formal query template. Consider a KG represented in RDF [^2] https://www.w3.org/RDF/ format and a corresponding SPARQL [^3] https://www.w3.org/TR/sparql11-query/ query, the relation chain elucidated in Section 3.2 should be represented as follows,
PREFIX ent: <http://www.kg/entity/>
PREFIX rel: <http://www.kg/relation/>
SELECT DISTINCT ?id ?label WHERE {
   ent:E0 rel:R0 ?x.
   ?x rel:R1 ?y.
   ?y rel:R2 ?z.
   ?z rel:R3 ?id.
   ?id rdfs:label ?label.
LIMIT 1
In this context, "ent" and "rel" serve as prefixes for entity and relation, respectively. The identifier "E0" uniquely represents "Eeyore" within the KG, while the identifiers for "creator," "child," "country of citizenship," and "capital" are denoted as "R0", "R1", "R2", and "R3" respectively. After identifying the entity "?id", we retrieve its string label "?label" as the final answer.
3.4 Integrating LLM-based QA and KBQA
This subsection outlines the integration of the proposed KBQA module with the LLM-based QA module within the GMeLLo framework.
LLM-based question answering. When a question arises, we retrieve the top-$x$ relevant facts using the pre-trained Contriever Izacard et al. (2022) model from a list of edited fact sentences. We then prompt the LLMs to generate answers based on the question and these pertinent facts. Compared to the "split-answer-check" pipeline in MeLLo Zhong et al. (2023), this LLM-based QA method is expected to be simpler and yield more accurate results when the facts are provided accurately.
However, addressing multi-hop questions, especially those where the edited facts pertain to intermediary hops, presents a challenge in accurately retrieving the relevant information and performing correct multi-hop question answering. This challenge is particularly pronounced when dealing with a large volume of edited facts. For instance, accurately identifying the relevant fact given the question in Figure  2 and producing the correct final answer is difficult.
KBQA. To address the challenges of LLM-based question answering, we integrate responses from KBQA to refine the outputs from the LLMs, as detailed in Sections 3.1-3.3. When the relation chain and fact triples are accurately derived, the KBQA system provides the correct answer. However, if the relation chain is incorrectly extracted, the search path in the KG may become invalid, leading the KBQA system to yield no output. In such instances, we accept the response from the LLMs as the final answer.
4 Experiment
In the upcoming section, we will conduct experiments to demonstrate the effectiveness of employing our GMeLLo methodology.
4.1 Experiment Setup
4.1.1 Dataset
Our experiment focuses on the multi-hop question-answering benchmark, MQuAKE Zhong et al. (2023), which comprises two datasets: MQuAKE-CF  [^4] Following Zhong et al. (2023), our experiments on MQuAKE-CF are carried out on a randomly sampled subset of the complete dataset, comprising 3000 instances in total(1000 instances for each of 2, 3, 4-hop questions)., designed for counterfactual edits, and MQuAKE-T, specifically tailored for updates in temporal knowledge.
The MQuAKE-CF dataset comprises 3,000 N-hop questions ($\textrm{N}\in\{2,3,4\}$), each linked to one or more edits. This dataset functions as a diagnostic tool for examining the effectiveness of knowledge editing methods in handling counterfactual edits. The MQuAKE-T dataset consists of 1,868 instances, each associated with a real-world fact change. Its purpose is to evaluate the efficacy of knowledge editing methods in updating obsolete information with contemporary, factual data. A table of statistics is available in Appendix A.1.
4.1.2 Evaluation Settings
To evaluate our models, we adhere to the testing settings outlined by Zhong et al. (2023). Specifically, instances are batched in groups of size $k$, with $k\in{1,100,1000,3000}$ for MQuAKE-CF, and $k\in{1,100,500,1868}$ for MQuAKE-T. For example, in the MQuAKE-CF dataset, when $k=100$, the $3000$ instances are split into $30$ groups, and we report the average performance as the final result.
For each test instance, the dataset includes three multi-hop questions that convey the same meaning. In alignment with Zhong et al. (2023), if the model correctly answers any one of these questions, we consider the instance to be accurately resolved.
4.1.3 Baselines
To demonstrate the effectiveness of our approach, we conduct comparisons with the following SOTA knowledge editing methods.
  • MEND Mitchell et al. (2022a). It trains a hyper-network to generate weight updates by transforming raw fine-tuning gradients based on an edited fact.
  • MEMIT Meng et al. (2023). It updates feed-forward networks across various layers to incorporate all relevant facts.
  • MeLLo Zhong et al. (2023). It employs a memory-based approach for multi-hop question answering, storing all updated facts in an external memory.
Given the substantial costs associated with training, deploying, and maintaining larger LLMs Li et al. (2023b), and the challenges of scaling up knowledge editing methods that require model parameter modifications, this paper primarily focuses on smaller LLMs, specifically GPT-J (6B) Wang and Komatsuzaki (2021) and Vicuna (7B) Chiang et al. (2023). However, to showcase GMeLLo’s effectiveness with larger LLMs in practical scenarios, we also report the performance of both MeLLo and GMeLLo on the MQuAKE-CF dataset when $k=3000$.
Base Model Method MQuAKE-CF MQuAKE-T
k=1 k=100 k=1000 k=3000 k=1 k=100 k=500 k=1868
GPT-J-6B MEMIT 12.3 9.8 8.1 1.8 4.8 1.0 0.2 0.0
MEND 11.5 9.1 4.3 3.5 38.2 17.4 12.7 4.6
MeLLo 20.3 12.5 10.4 9.8 85.9 45.7 33.8 30.7
GMeLLo 76.3 53.4 49.5 49.0 86.9 82.1 81.5 81.5
Vicuna-7B MeLLo 20.3 11.9 11.0 10.2 84.4 56.3 52.6 51.3
GMeLLo 71.3 46.5 42.5 41.9 97.1 86.3 85.4 85.1
Table 1: Performance comparison of GMeLLo and other approaches on the MQuAKE-CF and MQuAKE-T datasets using GPT-J-6B or Vicuna-7B as the base language models. Adhering to the methodology outlined by Zhong et al. (2023), instances are grouped into batches of size $k$. For the MQuAKE-CF dataset, $k$ varies from 1 to 3000, and for the MQuAKE-T dataset, it ranges from 1 to 1868. For example, in the MQuAKE-CF dataset, when $k=100$, the 3000 instances are organized into 30 groups, and the average performance reported as the final result. The metric used is accuracy.
4.1.4 Knowledge Graph Setting
Considering Wikidata’s community-driven nature, guaranteeing a dynamic and comprehensive dataset across a spectrum of knowledge domains, we use Wikidata Vrandečić and Krötzsch (2014) as the foundational KG for this experiment. To align the relations in the question and fact sentences with those in WikiData Vrandečić and Krötzsch (2014), we follow the following steps:
  • We select the first 500 item properties [^5] https://www.wikidata.org/w/index.php?title=Special:ListProperties/wikibase-item&limit=500&offset=0 from WikiData as the base relations. Items represent either concrete or abstract entities, such as a person Piscopo and Simperl (2019).
  • Next, we employ GPT-3.5-Turbo to examine each multi-hop question in the test samples to determine if it contains any of the base relations.
  • Afterward, we rank the frequencies of each relation and choose the top 50 relations as candidates for use in relation chain extraction and edited fact triple extraction.
To stay updated with the latest information on WikiData, we utilize the WikiData API service [^6] https://www.wikidata.org/w/api.php and the WikiData Query Service [^7] https://query.wikidata.org/sparql. The correctness of our KBQA result hinges on the accurate extraction of both edited fact triples and relation chains. If the relation chain is found to be incorrect, we conduct an online search on WikiData to determine if the relation chain leads to an entity that could potentially yield an incorrect answer for the specific question, which takes about 1 second.
4.1.5 Strategies for Managing Unforeseen Relationships
As previously noted, since LLMs may produce relations that are similar in meaning but not identical, we employ the pretrained Contriever model Izacard et al. (2022) to retrieve the most similar relation (i.e., the closest relation in the embedding space) from the base list of relations. This replacement is performed when undefined relations are encountered during both edited fact triple extraction and relation chain extraction.
4.2 Main Results
As shown in Table 1, our GMeLLo significantly outperforms all existing methods on the both the MQuAKE-CF dataset and the MQuAKE-T dataset Zhong et al. (2023), particularly when handling a large number of edits.
The performance degradation in MeLLo is primarily due to its challenges in identifying relevant facts as the number of edits increases. When k=1, the model utilizes only the facts directly related to the input question for context. However, as k increases, the model faces the challenge of discerning relevant facts from a broader memory. Our proposed GMeLLo model mitigates this by employing an explicit symbolic graph representation, which enhances the system’s ability to update and retrieve relevant facts effectively. This feature significantly boosts the scalability of GMeLLo, making it well-suited for real-world question answering applications that require managing large volumes of rapidly changing information.
In addition, we evaluated MeLLo and GMeLLo using two larger models, GPT-3.5-Turbo-Instruct and GPT-3.5-Turbo [^8] https://platform.openai.com/docs/models/gpt-3-5-turbo, on the MQuAKE-CF dataset with k=3000999The model text-davinci-003 used in Zhong et al. (2023) was deprecated on January 4, 2024.. The accuracy rates achieved by MeLLo and GMeLLo with GPT-3.5-Turbo-Instruct were 30.7% and 51.4%, respectively. While GMeLLo achieved an accuracy of 66.4% with GPT-3.5-Turbo, the same model consistently returned errors when tested with MeLLo, suggesting that the prompts may require modification for compatibility with chat completion models. These results indicate that GMeLLo performs well even when scaled to larger LLMs.
Base Model Method MQuAKE-CF MQuAKE-T
k=1 100 1000 3000 k=1 100 500 1868
GPT-J-6B QA 71.0 24.2 14.3 12.2 32.3 18.0 15.7 15.5
KBQA 43.3 43.3 43.3 43.3 80.2 80.2 80.2 80.2
GMeLLo 76.3 53.4 49.5 49.0 86.9 82.1 81.5 81.5
Vicuna-7B QA 72.6 27.0 16.5 13.5 96.9 63.0 59.2 58.2
KBQA 35.9 35.9 35.9 35.9 73.6 73.6 73.6 73.6
GMeLLo 71.3 46.5 42.5 41.9 97.1 86.3 85.4 85.1
Table 2: Ablation study of GMeLLo. QA involves directly using LLM for answering the multi-hop questions. KBQA involves using LLM to transform edited fact sentences into triples, update WikiData, convert question sentences into relation chains, and generate formal KG queries for question answering. GMeLLo combines these methods by using KBQA to correct answers from LLM-based QA.
4.3 Ablation Study
To gain a comprehensive understanding of the performance of various components, i.e., LLM-based QA and KBQA, we conduct an experiment to illustrate the impact of LLM-based QA and KBQA as the number of edits increases.
As demonstrated in Table  2, the performance of KBQA remains consistent, as all edited facts are converted to triples and all relation chains are extracted from the test questions, regardless of the value of ’k’. However, as the parameter ’k’ increases, more edited facts are stored in the external memory. Consequently, selecting the relevant edits to accurately answering the questions becomes increasingly challenging for LLM-based QA.
When k=1 and all relevant facts are provided to the LLMs for question answering, the LLM-based QA proves to be quite effective. However, a more realistic scenario involves multiple edits occurring simultaneously, where each question is asked separately (i.e., k>1). The performance showcased in Table 2 demonstrates the effectiveness of our GMeLLo, highlighting that KBQA serves as a valuable enhancement to LLM-based QA within evolving environments.
4.3.1 Further Analysis
To evaluate the impact of KBQA on LLM-based QA within the GMeLLo framework, we conducted an analysis comparing the responses from LLMs to those from the KG. We consider the KG’s response as the final answer. Therefore, comparing to only using LLM-based QA, if the answer from LLMs is correct but the answer from the KG is incorrect, this leads to a decline in performance. Conversely, if the answer from LLMs is incorrect but the answer from the KG is correct, performance improves. If the KBQA provides no response, performance remains unchanged. As illustrated in Table 3, when there are discrepancies between KBQA and LLM-based QA responses, the likelihood of KBQA providing the correct answer increases as the parameter k increases.
Base Model Scenario MQuAKE-CF MQuAKE-T
LLM KG Performance k=1 100 1000 3000 k=1 100 500 1868
GPT-J-6B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 8.1 22.9 24.9 25.0 44.0 47.2 47.9 48.0
${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 12.5 2.4 1.2 0.7 0.7 0.4 0.3 0.3
[…]
Vicuna-7B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 7.7 17.8 19.6 20.0 4.2 19.7 21.4 21.7
${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 21.8 3.9 2.0 1.2 7.2 4.2 4.0 3.9
[…]
4.4 Qualitative Analysis
Table  2 illustrates that Vicuna exhibits superior performance in directly handling the QA task, particularly when provided with the exact edited facts. Conversely, GPT-J excels in sentence analysis tasks, showcasing its high performance in the KBQA task.
4.4.1 Inferior Performance of GPT-J in QA
Table 2 shows that the performance of GPT-J and Vicuna in conducting QA tasks is comparable on the MQuAKE-CF dataset when k=1. However, GPT-J exhibits notably lower performance on the MQuAKE-T dataset. Further analysis revealed that GPT-J struggles in answering questions with only an edited fact pertaining to its intermediary information, such as: {examplebox} Sample from MQuAKE-CF
Facts: Midfielder is associated with the sport of Gaelic football
Question: What is the capital of the country where the sport associated with Kieron Dyer’s specialty was first played?
Predicted Answer: Bondi Junction
Answer: Dublin
{examplebox}
Sample from MQuAKE-T
Facts: The name of the current head of the Philippines government is Bongbong Marcos
Question: Who is the head of government of the country that Joey de Leon is a citizen of?
Predicted Answer: Benigno Aquino III
Answer: Bongbong Marcos
However, it can achieve the correct answer in KBQA because it accurately extracts the fact triple and relation chain of the question. Given that all test samples in MQuAKE-T contain only one edited fact, while approximately 63.6% of test samples in MQuAKE-CF consist of more than two edited facts, GPT-J is able to connect most of the information together. Therefore, it achieves better performance in the MQuAKE-CF dataset.
4.4.2 Inferior Performance of Vicuna in KBQA
Compared to GPT-J, Vicuna performs less effectively in the KBQA task. Aside from misunderstandings, the main reasons are as follows:
  • It often makes errors in the sequence. For example, given the fact "The author of Misery is Richard Dawkins", its output fact triple is "Richard Dawkins->author->Misery". However, the correct sequence is "Misery->author->Richard Dawkins".
  • It frequently makes errors in selecting a relation from the list. For example, it often outputs a relation chain as "Mike->citizenship->country->head of state", instead of "Mike->country of citizenship->head of state".
It is important to note that even if the relation chain is incorrect, the KBQA system may still provide the correct answer because of some loops in WikiData, such as the country of the USA is the USA.
Although Vicuna is not as effective overall, we still find that in some cases it can correctly extract relations, but cannot provide the correct answer directly. An example is given as follows:
{examplebox}
Sample from MQuAKE-CF
Facts: Point guard is associated with the sport of cricket
Question: What is the capital of the country from which Erik Spoelstra’s sport comes?
Predicted Answer: Miami
Answer: London
4.5 Further Discussion
KG offers a clearer representation of multi-hop information and its updates. In GMeLLo, we harness the strengths of both KBQA and LLM-based QA, benefiting from KBQA’s high precision and LLM-based QA’s extensive coverage. Our experiments reveal that GPT-J excels in extracting relation chains and fact triples, whereas Vicuna demonstrates superior performance in LLM-based QA. Given that KBQA and LLM-based QA operate as separate modules in GMeLLo, we can optimize their use by employing different LLMs in each module, maximizing their effectiveness in practical applications.
5 Conclusion
In this paper, we present GMeLLo, a method designed for multi-hop question answering in dynamic environments. Except leveraging LLMs for question answering, we also leverage the capabilities of LLMs to extract the triples from edited fact sentence to update KG, and use the capabilities of LLMs to analyze question sentences and generate a relation chain, and finally get the formal query by filling in a formal query template. Finally, we combine KBQA and LLM-based QA to bolster the multi-hop question answering capability within a dynamic environment. This approach capitalizes on the strengths of both LLMs and KGs. By utilizing LLMs for analyzing question sentences and QA to ensure the coverage, and KBQA to provide accurate results, we achieve a synergy between these two methodologies.
Limitations
Despite the promising results, it is important to acknowledge that this investigation is still in its early stages. Although our performance significantly surpasses baseline approaches in multi-hop questions in dynamic domains, particularly for large knowledge bases and complex questions, there is still room for further improvement. Our future research includes
  • Leveraging more sophisticated prompting techniques, such as Chain of Thought (CoT) Wei et al. (2022), to enable more accurate multi-hop reasoning.
  • Refining the predefined relation list to enhance its accuracy.
  • Enhancing the KG to support more complex question answering, such as inquiries involving historical information.
We believe these improvements can further enhance the performance and scalability of the system, enabling it to handle more complex and diverse real-world applications.
Appendix A Implementation
A.1 Dataset Statistics
Table 4 provides a summary of the statistics for the MQuAKE-CF and MQuAKE-T datasets.
#Edits 2-hop 3-hop 4-hop Total
MQuaKE-CF 1 513 356 224 1,093
2 487 334 246 1,067
3 - 310 262 572
4 - - 268 268
All 1,000 1,000 1,000 3,000
MQuaKE-T 1 (All) 1,421 445 2 1,868
Table 4: Statistics of MQuAKE dataset Zhong et al. (2023).
k=100 k=1000 k=3000 Average
Top-4 15.6 9.1 7.2 10.63
Top-5 16.8 8.3 6.9 10.67
Top-6 16.6 8.5 7.4 10.83
Top-10 15.3 9.0 8.0 10.77
Top-100 8.2 4.7 3.7 5.53
Table 5: Hyperparameter search for top-$x$ in Vicuna-based QA systems on the MQuAKE-CF dataset.
A.2 Hyperparmers Settings
To ensure reproducibility, we set the temperature to zero in all experiments. Table 5 shows that retrieving the top-6 edited facts from external memory provides the best average performance on the MQuAKE-CF dataset for $k>1$. Consequently, we include top-6 edited facts in the prompt for subsequent experiments on this dataset when $k>1$. Similarly, for the MQuAKE-T dataset when $k>1$, we opted to incorporate the top-1 edited fact in the prompt.
A.3 Predefined Relations Utilized in the Prompts for Relation Chain and Fact Triple Extraction
After filtering by GPT-3.5-Turbo, the first 50 relations utilized in MQuAKE-CF dataset are: [’country of origin’,’sport’, ’country of citizenship’, ’capital’, ’continent’, ’official language’, ’head of state’, ’head of government’, ’creator’, ’country’, ’author’, ’headquarters location’, ’place of birth’,’spouse’, ’director / manager’,’religion or worldview’, ’genre’, ’work location’, ’performer’,’manufacturer’, ’developer’, ’place of death’, ’employer’, ’educated at’,’member of sports team’, ’head coach’, ’languages spoken, written or signed’, ’notable work’, ’child’, ’founded by’, ’location’, ’chief executive officer’, ’original broadcaster’, ’chairperson’, ’occupation’, ’position played on team / speciality’,’member of’, ’language of work or name’, ’director’, ’league’, ’home venue’, ’native language’, ’composer’, ’place of origin (Switzerland)’, ’officeholder’,’religious order’, ’publisher’, ’original language of film or TV show’, ’ethnic group’,’military branch’].
After GPT-3.5-Turbo filtering, the MQuAKE-T dataset includes a total of 35 relations. The relation list is [’head of government’, ’country of citizenship’, ’head of state’, ’country of origin’, ’country’, ’headquarters location’, ’location’, ’sport’, ’performer’, ’genre’, ’developer’, ’employer’, ’manufacturer’, ’place of death’, ’place of birth’, ’author’, ’member of’, ’capital’, ’member of sports team’, ’chief executive officer’, ’notable work’, ’director / manager’, ’original broadcaster’, ’creator’, ’work location’, ’educated at’, ’located in the administrative territorial entity’, ’head coach’, ’place of publication’, ’location of formation’, ’director’, ’producer’, ’transport network’, ’continent’, ’child’]
A.4 Prompt Setup and Post-Processing
The prompts used for edited fact triple extraction, relation chain extraction, and LLM-based QA are depicted in Figures  3,  4, and  5. The edited triple can be regarded as a specialized relation chain, with only one relation between entities and all entities known. All samples in the prompt are selected from the complete MQuAKE-CF dataset, ensuring they are distinct from the test samples.
Figure 3: The prompt used for transforming edited fact sentences to triples.
Figure 4: The prompt used for transforming question sentences to relation chains.
Figure 5: The prompt used in LLM-based QA.
To improve the performance of LLMs in extracting relation chains and ensure that outputs conform to a specified format, we employ a 4-shot learning approach for the MQuAKE-CF dataset and a 3-shot learning approach for the MQuAKE-T dataset. For MQuAKE-CF, the approach involves presenting the model with samples of one 2-hop question, one 3-hop question, and two 4-hop questions. In contrast, for MQuAKE-T, the model is presented with one 2-hop question, one 3-hop question, and one 4-hop question.
To address the limitations of GPT-J and Vicuna in conforming to the desired output format, we establish a heuristic rule for extracting essential information from their outputs. For instance, in the context of relation chain extraction, this heuristic is outlined as follows:
  • Narrow the attention to the output sentence containing the "->" indicator.
  • Divide the sentence based on the "->" delimiter.
  • Regard the initial segment as the predicted entity. Subsequently, process the following segments sequentially as relations, provided they do not begin with "?".
A.5 Strategies for Managing Sequence Errors in Extracting Fact Triples
While LLMs consistently identifies relations accurately—such as ’head of state,’ ’chief of department,’ and ’head of government’—it often makes errors in their sequencing. To address this, we employ Spacy [^10] https://spacy.io/ to detect instances where the object of an edited triple is not a person. If it is not, we adjust the sequence of the object and subject in the triple accordingly.
Appendix B The Distinctions Between Our GMeLLo and Other Methods
While both GMeLLo and MeLLo Zhong et al. (2023) are memory-based models targeting multi-hop question answering in an evolving environment, they differ in the following aspects:
  • MeLLo employs in-context learning to direct LLMs in splitting the question into sub-questions, answering each, and verifying against relevant edited facts for contradictions. In contrast, GMeLLo retrieves pertinent edited facts for the multi-hop question and presents them alongside the question to LLMs for answering.
  • Except storing edited facts as isolated sentences in an external memory, we leverage LLMs to translate these sentences into triples and update the KG. In addition to obtaining an answer from LLMs, we utilize KBQA to enhance the precision of multi-hop question answering within an evolving environment.
Recently, the advent of LLMs has spurred the development of LLM-based KBQA systems Baek et al. (2023); Sen et al. (2023); Nie et al. (2024). However, our GMeLLo are different from these works in the following aspects:
  • Firstly, we consider question answering in a dynamic environment, where changes in the knowledge graph need to accounted for, whereas they do not.
  • Secondly, we focus on multi-hop questions, whereas they deal with standard KBQA tasks, including intersection and difference questions etc.
  • Thirdly, the KBQA and LLM-based QA are handled separately, using the KBQA answer as the final answer. In contrast, they retrieve triples from the knowledge graph and incorporate them into the prompt to guide LLM-based QA.
Appendix C Multi-Hop Performance Analysis
We study the breakdown of performance on the MQuAKE-CF dataset with respect to the number of hops when $k=1$. Table 6 provides the hop-specific performance of different methods. Although MQuAKE did not provide the hop performance for MeLLo, it can be inferred that the average hop performance should not exceed 65.6%, given that the overall performance is 20.3%.
Model Method Number of Hops
2 3 4 Avg
GPT-J-6B MEND 13.9 11.3 9.5 11.5
MEMIT 22.5 6.0 8.4 12.3
MeLLo - - - 20.3
GMeLLo 89.5 73.7 65.6 76.3
Table 6: The breakdown performance on the MQuAKE-CF dataset with respect to the number of hops when $k=1$.

### v2

@@
 Abstract
-The rapid obsolescence of information in Large Language Models (LLMs) has driven the development of various techniques to incorporate new facts.
-However, existing methods for knowledge editing still face difficulties with multi-hop questions that require accurate fact identification and sequential logical reasoning, particularly among numerous fact updates.
-To tackle these challenges, this paper introduces Graph Memory-based Editing for Large Language Models (GMeLLo), a straitforward and effective method that merges the explicit knowledge representation of Knowledge Graphs (KGs) with the linguistic flexibility of LLMs.
+The important challenge of keeping knowledge in Large Language Models (LLMs) up-to-date has led to the development of various methods for incorporating new facts.
+However, existing methods for such knowledge editing still face difficulties with multi-hop questions that require accurate fact identification and sequential logical reasoning, particularly among numerous fact updates.
+To tackle these challenges, this paper introduces Graph Memory-based Editing for Large Language Models (GMeLLo), a straightforward and effective method that merges the explicit knowledge representation of Knowledge Graphs (KGs) with the linguistic flexibility of LLMs.
 Beyond merely leveraging LLMs for question answering, GMeLLo employs these models to convert free-form language into structured queries and fact triples, facilitating seamless interaction with KGs for rapid updates and precise multi-hop reasoning.
-Our results show that GMeLLo significantly surpasses current state-of-the-art knowledge editing methods in the multi-hop question answering benchmark, MQuAKE, especially in scenarios with extensive knowledge edits.
+Our results show that GMeLLo significantly surpasses current state-of-the-art (SOTA) knowledge editing methods in the multi-hop question answering benchmark, MQuAKE, especially in scenarios with extensive knowledge edits.
 […]
@@
 For instance, a transition in the British Prime Minister, such as from Boris Johnson to Rishi Sunak, necessitates corresponding adjustments, like the change in the British Prime Minister’s spouse.
-As the widespread deployment of LLMs continues, the imperative to keep their knowledge accurate and up-to-date, without incurring extensive retraining costs, becomes increasingly evident Sinitsin et al.
+An important challenge in deploying Large Language Models (LLMs) is keeping their knowledge accurate and up-to-date, without incurring expensive retraining costs Sinitsin et al.
 (2020).
-Several approaches have been proposed in prior works to address this challenge, with some focusing on the incremental injection of new facts into language models Rawat et al.
+Several approaches have been proposed in prior works to address this challenge.
+Some methods focus on the incremental injection of new facts into language models Rawat et al.
 (2020); De Cao et al.
@@
 (2022a).
-Interestingly, certain methodologies in the literature diverge from the conventional path of updating model weights, opting instead for an innovative strategy involving the use of external memory to store the edits Mitchell et al.
+Alternatively, other methods involve the use of external memory to store new facts Mitchell et al.
 (2022b); Zhong et al.
+(2023), which does not require updating LLM model weights.
+As LLMs operate as black boxes, modifying one fact might inadvertently alter another, making it challenging to guarantee accurate revisions.
+In this paper, we introduce GMeLLo, an effective approach designed to synergize the strengths of LLMs and Knowledge Graphs (KGs) in addressing the multi-hop question answering task after knowledge editing Zhong et al.
 (2023).
-As LLMs operate as black boxes, modifying one fact might inadvertently alter another, making it challenging to guarantee accurate revisions.
-In this paper, we introduce GMeLLo, an effective approach designed to synergize the strengths of LLMs and KGs in addressing the multi-hop question answering task after knowledge editing Zhong et al.
-(2023).
-An illustrative example of our focus is presented in Figure  1.
-Following an update regarding the information of the British Prime Minister, it becomes evident that the corresponding spouse information should also be modified.
-As depicted in Figure  2, our GMeLLo comprises the following key steps:
+An illustrative example is presented in Figure 1.
+Following an information update regarding the British Prime Minister, it becomes evident that the corresponding spouse information should also be modified.
+As depicted in Figure 2, our GMeLLo method comprises the following key steps:
   • We utilize LLMs to translate edited fact sentences into triples, employing these triples to update the KG and ensure its information remains up to date.
@@
 In cases where KBQA fails, we still have LLMs for question answering (QA) to ensure comprehensive coverage.
-GMeLLo outperforms current state-of-the-art (SOTA) methods on two datasets from the MQuAKE benchmark, affirming its effectiveness in multi-hop question answering within an evolving environment.
+GMeLLo outperforms current SOTA methods on two datasets from the MQuAKE benchmark, affirming its effectiveness in multi-hop question answering within an evolving environment.
 2 Related Work
@@
 As highlighted in Yao et al. (2023), two paradigms exist for editing knowledge: modifying model parameters and preserving model parameters.
-2.2.1 Modifying Model Parameters
+2.2.1 Parameter-Modification Paradigm
 In the case of modifying model parameters, this can be further categorized into meta-learning or locate-and-edit approaches.
-Meta-learning methods, as discussed in De Cao et al.
+Meta-learning methods De Cao et al.
 (2021); Mitchell et al.
-(2022a), utilize a hyper network to learn the necessary adjustments for editing LLMs.
-The locate-then-edit paradigm, as demonstrated in Dai et al.
+(2022a) utilize a hyper network to learn the necessary adjustments for editing LLMs.
+The locate-then-edit paradigm Dai et al.
 (2022); Meng et al.
@@
 (2023); Zhang et al.
-(2024), involves initially identifying parameters corresponding to specific knowledge and subsequently modifying them through direct updates to the target parameters.
-2.2.2 Preserving Model Parameters
+(2024) involves initially identifying parameters corresponding to specific knowledge and subsequently modifying them through direct updates to the target parameters.
+2.2.2 Parameter-Preservation Paradigm
 In the case of preserving model parameters, the introduction of additional parameters or external memory becomes necessary.
-The paradigm of additional parameters, as presented in Dong et al.
+The paradigm of additional parameters Dong et al.
 (2022); Hartvigsen et al.
 (2022); Huang et al.
-(2022), incorporates extra trainable parameters into the language model.
+(2022) incorporates extra trainable parameters into the language model.
 These parameters are trained on a modified knowledge dataset, while the original model parameters remain static.
-On the other hand, memory-based models Mitchell et al.
+In contrast, memory-based models Mitchell et al.
 (2022b); Zhong et al.
-(2023) explicitly store all edited examples in memory and employ a retriever to extract the relevant edit facts for each new input, guiding the model in generating the edited output.
+(2023); Gu et al.
+(2024) explicitly store all edited examples in memory and employ a retriever to extract the relevant edit facts for each new input, guiding the model in generating the updated output.
 While previous evaluation paradigms have primarily focused on validating the recall of edited facts, Zhong et al.
@@
 When updated facts are received, we utilize LLMs to extract entities from the sentences and determine their relationships (selecting a relation from the predefined list).
-This process generates edited fact triples, which are then used to update the KG (see Figure  2).
+This process generates edited fact triples, which are then used to update the KG (see Figure 2).
 Updating the KG with an edited fact triple involves identifying the connections in the KG based on the subject entity and relation, breaking these connections, and establishing a new connection based on the triple.
@@
 In this paper, we employ LLMs to extract the relation chain from a sentence, encompassing the mentioned entity in the question and its relations with other unidentified entities.
-Similar to the fact triple exaction 3.1, we task LLMs with selecting a relation from a predefined list to mitigate varied representations of the same relation.
+Similar to the fact triple exaction mentioned in Section 3.1, we task LLMs with selecting a relation from a predefined list to mitigate varied representations of the same relation.
 Take a question sentence from the MQuAKE-CF Zhong et al.
@@
 The presented question necessitates a 4-hop reasoning process.
-With "Eeyore" as the known entity in focus, the journey to the final answer involves identifying its creator ’?x’, moving on to the creator’s child ’?y’, obtaining the child’s country of citizenship ’?z’, and culminating with the retrieval of the country’s capital ’?m’.
-All the relations, such as ’creator,’ ’child,’ ’country of citizenship,’ and ’capital,’ are chosen from a predefined list of relations.
+With "Eeyore" as the known entity in focus, the journey to the final answer involves identifying its creator "?x", moving on to the creator’s child "?y", obtaining the child’s country of citizenship "?z", and culminating with the retrieval of the country’s capital "?m".
+All the relations, such as "creator", "child", "country of citizenship", and "capital", are chosen from a predefined list of relations.
 The relation chain encapsulates all essential information for deriving the answer.
@@
 (2023), as detailed in Appendix A.4.
-3.3 Converting Relation Chain into a Formal Query
+3.3 Converting a Relation Chain into a Formal Query
 Once the relation chain is obtained, the next step involves integrating the known entity and the relations into a formal query template.
-Consider a KG represented in RDF [^2] https://www.w3.org/RDF/ format and a corresponding SPARQL [^3] https://www.w3.org/TR/sparql11-query/ query, the relation chain elucidated in Section 3.2 should be represented as follows,
+For a KG represented in RDF [^2] https://www.w3.org/RDF/ format, the relation chain elucidated in Section 3.2 can be represented as the following SPARQL [^3] https://www.w3.org/TR/sparql11-query/ query,
 PREFIX ent: <http://www.kg/entity/>
@@
 In this context, "ent" and "rel" serve as prefixes for entity and relation, respectively.
-The identifier "E0" uniquely represents "Eeyore" within the KG, while the identifiers for "creator," "child," "country of citizenship," and "capital" are denoted as "R0", "R1", "R2", and "R3" respectively.
+The identifier "E0" uniquely represents "Eeyore" within the KG, while the identifiers for "creator," "child," "country of citizenship," and "capital" are denoted as "R0", "R1", "R2", and "R3", respectively.
 After identifying the entity "?id", we retrieve its string label "?label" as the final answer.
@@
 This challenge is particularly pronounced when dealing with a large volume of edited facts.
-For instance, accurately identifying the relevant fact given the question in Figure  2 and producing the correct final answer is difficult.
+For instance, accurately identifying the relevant fact given the question in Figure 2 and producing the correct final answer is difficult.
 KBQA.
-To address the challenges of LLM-based question answering, we integrate responses from KBQA to refine the outputs from the LLMs, as detailed in Sections 3.1-3.3.
+To address the challenges of LLM-based question answering, we integrate responses from KBQA to refine the outputs from the LLMs, as detailed in the previous section.
 When the relation chain and fact triples are accurately derived, the KBQA system provides the correct answer.
@@
 4 Experiment
-In the upcoming section, we will conduct experiments to demonstrate the effectiveness of employing our GMeLLo methodology.
+In this section, we will present the results from our experiments to demonstrate the effectiveness of employing our GMeLLo methodology.
 4.1 Experiment Setup
@@
   • MeLLo Zhong et al. (2023). It employs a memory-based approach for multi-hop question answering, storing all updated facts in an external memory.
+  • PokeMQA Gu et al.
+(2024).
+It also uses a memory-based approach, which decouples question decomposition from knowledge editing to reduce the burden on LLMs.
+Additionally, it introduces auxiliary knowledge prompts to assist with question decomposition.
 Given the substantial costs associated with training, deploying, and maintaining larger LLMs Li et al.
@@
 Vicuna-7B MeLLo 20.3 11.9 11.0 10.2 84.4 56.3 52.6 51.3
+PokeMQA 45.8 38.8 - 31.6 74.6 - - 73.1
 GMeLLo 71.3 46.5 42.5 41.9 97.1 86.3 85.4 85.1
@@
 Considering Wikidata’s community-driven nature, guaranteeing a dynamic and comprehensive dataset across a spectrum of knowledge domains, we use Wikidata Vrandečić and Krötzsch (2014) as the foundational KG for this experiment.
-To align the relations in the question and fact sentences with those in WikiData Vrandečić and Krötzsch (2014), we follow the following steps:
-  • We select the first 500 item properties [^5] https://www.wikidata.org/w/index.php?title=Special:ListProperties/wikibase-item&limit=500&offset=0 from WikiData as the base relations.
-Items represent either concrete or abstract entities, such as a person Piscopo and Simperl (2019).
-  • Next, we employ GPT-3.5-Turbo to examine each multi-hop question in the test samples to determine if it contains any of the base relations.
-  • Afterward, we rank the frequencies of each relation and choose the top 50 relations as candidates for use in relation chain extraction and edited fact triple extraction.
+To align the relations in the question and fact sentences with those in WikiData Vrandečić and Krötzsch (2014), we take the following steps:
 To stay updated with the latest information on WikiData, we utilize the WikiData API service [^7] https://www.wikidata.org/w/api.php and the WikiData Query Service [^8] https://query.wikidata.org/sparql.
@@
 4.2 Main Results
-As shown in Table 1, our GMeLLo significantly outperforms all existing methods on the both the MQuAKE-CF dataset and the MQuAKE-T dataset Zhong et al.
+As shown in Table 1999We use the baseline performance reported in Zhong et al.
+(2023) and Gu et al.
+(2024).
+Since the experiment settings in Gu et al.
+(2024) differ from those in Zhong et al.
+(2023), we only include the results from Gu et al.
+(2024) under the same settings.
+A dash (’-’) indicates that performance was not reported for that setting., our GMeLLo significantly outperforms all existing methods on the both the MQuAKE-CF dataset and the MQuAKE-T dataset Zhong et al.
 (2023), particularly when handling a large number of edits.
-The performance degradation in MeLLo is primarily due to its challenges in identifying relevant facts as the number of edits increases.
-When k=1, the model utilizes only the facts directly related to the input question for context.
-However, as k increases, the model faces the challenge of discerning relevant facts from a broader memory.
-Our proposed GMeLLo model mitigates this by employing an explicit symbolic graph representation, which enhances the system’s ability to update and retrieve relevant facts effectively.
-This feature significantly boosts the scalability of GMeLLo, making it well-suited for real-world question answering applications that require managing large volumes of rapidly changing information.
-In addition, we evaluated MeLLo and GMeLLo using two larger models, GPT-3.5-Turbo-Instruct and GPT-3.5-Turbo [^8] https://platform.openai.com/docs/models/gpt-3-5-turbo, on the MQuAKE-CF dataset with k=3000999The model text-davinci-003 used in Zhong et al.
-(2023) was deprecated on January 4, 2024..
-The accuracy rates achieved by MeLLo and GMeLLo with GPT-3.5-Turbo-Instruct were 30.7% and 51.4%, respectively.
-While GMeLLo achieved an accuracy of 66.4% with GPT-3.5-Turbo, the same model consistently returned errors when tested with MeLLo, suggesting that the prompts may require modification for compatibility with chat completion models.
-These results indicate that GMeLLo performs well even when scaled to larger LLMs.
 Base Model Method MQuAKE-CF MQuAKE-T
@@
 GMeLLo combines these methods by using KBQA to correct answers from LLM-based QA.
+The performance degradation in MeLLo is primarily due to its challenges in identifying relevant facts as the number of edits increases.
+When k=1, the model utilizes only the facts directly related to the input question for context.
+However, as k increases, the model faces the challenge of discerning relevant facts from a broader memory.
+Our proposed GMeLLo model mitigates this by employing an explicit symbolic graph representation, which enhances the system’s ability to update and retrieve relevant facts effectively.
+This feature significantly boosts the scalability of GMeLLo, making it well-suited for real-world question answering applications that require managing large volumes of rapidly changing information.
+To further validate our findings with more capable base models, we evaluated MeLLo and GMeLLo using two larger models, GPT-3.5-Turbo-Instruct and GPT-3.5-Turbo, on the MQuAKE-CF dataset with k=3000101010The model text-davinci-003 used in Zhong et al.
+(2023) was deprecated on January 4, 2024..
+The accuracy rates achieved by MeLLo and GMeLLo with GPT-3.5-Turbo-Instruct were 30.7% and 51.4%, respectively.
+While GMeLLo achieved an accuracy of 66.4% with GPT-3.5-Turbo, the same model consistently returned errors when tested with MeLLo, suggesting that the prompts may require modification for compatibility with chat completion models.
+These results indicate that GMeLLo performs well even when scaled to larger LLMs.
 4.3 Ablation Study
 To gain a comprehensive understanding of the performance of various components, i.e., LLM-based QA and KBQA, we conduct an experiment to illustrate the impact of LLM-based QA and KBQA as the number of edits increases.
-As demonstrated in Table  2, the performance of KBQA remains consistent, as all edited facts are converted to triples and all relation chains are extracted from the test questions, regardless of the value of ’k’.
-However, as the parameter ’k’ increases, more edited facts are stored in the external memory.
+As demonstrated in Table 2, the performance of KBQA remains consistent because all edited facts are converted to triples and all relation chains are extracted from the test questions, regardless of the value of "k".
+Correctly answering a multi-hop question in KBQA requires both accurate extraction of fact triples and the relation chain.
+However, as the parameter "k" increases, more edited facts are stored in the external memory.
 Consequently, selecting the relevant edits to accurately answering the questions becomes increasingly challenging for LLM-based QA.
@@
 If the KBQA provides no response, performance remains unchanged.
-As illustrated in Table 3, when there are discrepancies between KBQA and LLM-based QA responses, the likelihood of KBQA providing the correct answer increases as the parameter k increases.
+As illustrated in Table 3, when there are discrepancies between KBQA and LLM-based QA responses, the likelihood of KBQA providing the correct answer increases as the parameter "k" increases.
 Base Model Scenario MQuAKE-CF MQuAKE-T
 LLM KG Performance k=1 100 1000 3000 k=1 100 500 1868
-GPT-J-6B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 8.1 22.9 24.9 25.0 44.0 47.2 47.9 48.0
-${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 12.5 2.4 1.2 0.7 0.7 0.4 0.3 0.3
+GPT-J-6B ${\color[rgb]{0,1,0}\boldsymbol{\uparrow}}{}$ 8.1 22.9 24.9 25.0 44.0 47.2 47.9 48.0
+${\color[rgb]{1,0,0}\boldsymbol{\downarrow}}{}$ 12.5 2.4 1.2 0.7 0.7 0.4 0.3 0.3
 […]
-Vicuna-7B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 7.7 17.8 19.6 20.0 4.2 19.7 21.4 21.7
-${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 21.8 3.9 2.0 1.2 7.2 4.2 4.0 3.9
+Vicuna-7B ${\color[rgb]{0,1,0}\boldsymbol{\uparrow}}{}$ 7.7 17.8 19.6 20.0 4.2 19.7 21.4 21.7
+${\color[rgb]{1,0,0}\boldsymbol{\downarrow}}{}$ 21.8 3.9 2.0 1.2 7.2 4.2 4.0 3.9
 […]
@@
 4.4 Qualitative Analysis
-Table  2 illustrates that Vicuna exhibits superior performance in directly handling the QA task, particularly when provided with the exact edited facts.
+Table 2 illustrates that Vicuna exhibits superior performance in directly handling the QA task, particularly when provided with the exact edited facts.
 Conversely, GPT-J excels in sentence analysis tasks, showcasing its high performance in the KBQA task.
@@
 However, GPT-J exhibits notably lower performance on the MQuAKE-T dataset.
-Further analysis revealed that GPT-J struggles in answering questions with only an edited fact pertaining to its intermediary information, such as: {examplebox} Sample from MQuAKE-CF
+Further analysis revealed that GPT-J struggles in answering questions with only an edited fact pertaining to its intermediary information, such as:
+{examplebox}
+Sample from MQuAKE-CF
 Facts: Midfielder is associated with the sport of Gaelic football
@@
 In this paper, we present GMeLLo, a method designed for multi-hop question answering in dynamic environments.
-Except leveraging LLMs for question answering, we also leverage the capabilities of LLMs to extract the triples from edited fact sentence to update KG, and use the capabilities of LLMs to analyze question sentences and generate a relation chain, and finally get the formal query by filling in a formal query template.
+In addition to leveraging LLMs for question answering, we also leverage the capabilities of LLMs to extract the triples from edited fact sentences to update a KG, and use the capabilities of LLMs to analyze question sentences and generate a relation chain, and finally get the formal query by filling in a formal query template.
 Finally, we combine KBQA and LLM-based QA to bolster the multi-hop question answering capability within a dynamic environment.
@@
 We believe these improvements can further enhance the performance and scalability of the system, enabling it to handle more complex and diverse real-world applications.
+Acknowledgments
+We thank all reviewers for providing valuable feedback.
+This work was partially supported by A*STAR CRF funding awarded to Cheston Tan.
+Bo Xiong is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-1574 – 471687386.
 Appendix A Implementation
@@
 To ensure reproducibility, we set the temperature to zero in all experiments.
-Table 5 shows that retrieving the top-6 edited facts from external memory provides the best average performance on the MQuAKE-CF dataset for $k>1$.
+Table 5 shows that retrieving the top-6 edited facts from external memory provides the best average performance on the MQuAKE-CF dataset [^11] Tested only on the first question of each test instance, rather than all three for $k>1$.
 Consequently, we include top-6 edited facts in the prompt for subsequent experiments on this dataset when $k>1$.
@@
 A.4 Prompt Setup and Post-Processing
-The prompts used for edited fact triple extraction, relation chain extraction, and LLM-based QA are depicted in Figures  3,  4, and  5.
+The prompts used for edited fact triple extraction, relation chain extraction, and LLM-based QA are depicted in Figures 3, 4, and 5.
 The edited triple can be regarded as a specialized relation chain, with only one relation between entities and all entities known.
@@
 For MQuAKE-CF, the approach involves presenting the model with samples of one 2-hop question, one 3-hop question, and two 4-hop questions.
-In contrast, for MQuAKE-T, the model is presented with one 2-hop question, one 3-hop question, and one 4-hop question.
+For MQuAKE-T, the model is presented with one 2-hop question, one 3-hop question, and one 4-hop question.
 To address the limitations of GPT-J and Vicuna in conforming to the desired output format, we establish a heuristic rule for extracting essential information from their outputs.